A sample of gas at 25.00oC has a pressure of 92.50kPa. What temperature would change its pressure to 110.0kPa?
1
Expert's answer
2020-10-15T03:24:48-0400

We will use the relationship;

"\dfrac{P1V1}{T1}=\dfrac{P2V2}{T2}"


From the question it seems Volume is kept constant. When that is the case, the relationship becomes;

"\dfrac{P1}{T1}=\dfrac{P2}{T2}" (at constant volume)


P1=92.50kPa

P2 = 110.0kPa

T1 = 25 + 273 = 298K

T2 = ?


Solving for T2, the expression changes to;


"T2 = \dfrac{P2 x T1}{P1}"


"T2 = \dfrac{110.0kPa x 298K}{95.50kPa} = 343.246K"


Changing to oC, T2 = 70.25 oC